OpenStudy (anonymous):

What is the distance between points (4, −7) and (−5, 9)? Round to the nearest tenth of a unit.

OpenStudy (anonymous):

ok subtract x from x and y from y

OpenStudy (anonymous):

4 - -5 and -7 - 9

OpenStudy (anonymous):

nm

OpenStudy (welshfella):

do you know the distance formula between 2 given points?

OpenStudy (anonymous):

d= \[\sqrt{(x^2-x^1)^2}+(y^2-y^1)^2\]

OpenStudy (welshfella):

if the 2 points are (x1,y1) and (x2,y2) distance = sqrt((y2-y1)^2 + (x2-x1)^2))

OpenStudy (anonymous):

-5 - 4 and ^2 them =)

OpenStudy (anonymous):

then get the sqrt of that answer =)

OpenStudy (anonymous):

3 ...

OpenStudy (welshfella):

plug in x1 = 4, x2 = -5 , y1 = -7 and y2 = 9 into the formula and work it out

OpenStudy (anonymous):

now do 3+(9- -7)^2

OpenStudy (anonymous):

9- -7 = 16

OpenStudy (anonymous):

16^2 = ?

OpenStudy (anonymous):

now you finish it up

OpenStudy (anonymous):

256

OpenStudy (anonymous):

????

OpenStudy (anonymous):

i screwed up lol

OpenStudy (anonymous):

oh well d =18.4

OpenStudy (anonymous):

d = distance

OpenStudy (anonymous):

thats the anwser

OpenStudy (anonymous):

...

OpenStudy (anonymous):

yes =)